OpenStudy (anonymous):

Find the general solution to the differential equation: d^2y/dx^2 - 2(dy/dx)+4y = 2+3x+sinx

OpenStudy (anonymous):

First you have to solve the auxiliary equation. That will be: \[\lambda^2-2\lambda+4=0\]

OpenStudy (anonymous):

This one is a nasty one as it does not have real roots, but that is not a problem.

OpenStudy (anonymous):

I have to go for 5 min now but I will continue later if you wish

OpenStudy (anonymous):

thanks mate for an assignment worth a few marks would be appreciated.

OpenStudy (anonymous):

So I'm back. Solving the above equation for lamdba needs complex numbers but I guess that you know that. Solutions are: \[1+i \sqrt{3}\] \[1-i \sqrt{3}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

If the auxiliary equation has complex roots alfa± i, then two linearly independent solutions to ay'' + by' + cy = 0 are e^alfa t cos t and e^alfa t sin t,

OpenStudy (anonymous):

So the general solution is: \[Ae^tcos \sqrt{3}t + Be^tsin \sqrt{3}t\]

OpenStudy (anonymous):

This is just the 0 part

OpenStudy (anonymous):

All the t-s should be x

OpenStudy (anonymous):

2+3x+sinx for this you can do it in two parts first for the 2+3x d^2y/dx^2 - 2(dy/dx)+4y -3/4 x^2+1/8

OpenStudy (anonymous):

If I am right but I need to double check it, just a sec my girlfriend is calling me.

OpenStudy (anonymous):

no I am stupid.

OpenStudy (anonymous):

no your not thanks you are on the right path

OpenStudy (anonymous):

it has to be in the form of ax+b than y''=0 y'=a y=ax+b -2a+4(ax+b)=2+3x 4ax-2a+4b=2+3x a=3/4 b=8/7

OpenStudy (anonymous):

So it will have a part: 3/4x+8/7 Can you check this while I make a tea :-)

OpenStudy (anonymous):

yeah no worries

OpenStudy (anonymous):

OK for the sinx bit y has to be in the form of y= Qsinx+Wcosx than y'=Qcosx-Wsinx y''=-Qsinx-Wcosx

OpenStudy (anonymous):

-Qsinx-Wcosx-2(Qcosx-Wsinx)+4(Qsinx+Wcosx)=sinx 3Wcosx-2Qcosx+3Qsinx+2Wsinx=sinx 3W=2Q 3Q+2W=1 Q=3/2 W 3(3/2 W)+2W=1 13/2W=1 W=2/13 Q=3/13

OpenStudy (anonymous):

And we are done. The solution is the general sol above + the linear bit+the sin/cos bit

OpenStudy (anonymous):

Happy? :-)

OpenStudy (anonymous):

A and B from the general solution are constants. If more info is given than you can calculate it

OpenStudy (anonymous):

Andras thankyou very much.

OpenStudy (anonymous):

just one thing is your 8/7 correct or should it be 7/8 ?

OpenStudy (anonymous):

ur welcome, hope you get a few points! What course is it?

OpenStudy (anonymous):

not sure, you should do the calculation too. On the computer I dont see it through so well.

OpenStudy (anonymous):

Engineering Maths III I will work through my notes and your answers. Thanks again

OpenStudy (anonymous):

University In Australia

OpenStudy (anonymous):

Andras if possible could you please explain this part, maybe show all the working, I keep getting stuck. I am not quite sure of the algebra. Thanks it has to be in the form of ax+b than y''=0 y'=a y=ax+b -2a+4(ax+b)=2+3x 4ax-2a+4b=2+3x a=3/4 b=8/7

OpenStudy (anonymous):

In the equation you have 2+3x This can only come from some kind of polynomial. If it has x^2 or higher power than that part cannot disappear. So the polynomial has to be ax+b (a and b are constants) Do you need help for solving it?

OpenStudy (anonymous):

Andras, thanks for your reply, I was able to work through it myself. Its all good now.

OpenStudy (anonymous):

OK, great!